OpenStudy (anonymous):

(4x^6)^3/2 simplify the expression

OpenStudy (jdoe0001):

hmmm got a typo there

OpenStudy (jdoe0001):

\(\Large{ (4x^6)^{\frac{3}{2}}\implies (4)^{\frac{3}{2}}(x^6)^{\frac{3}{2}}\implies (2^2)^{\frac{3}{2}}(x^6)^{\frac{3}{2}} \\ \quad \\ (2)^{\cancel{ 2}\frac{3}{\cancel{ 2}}}(x)^{\cancel{ 6}\cdot \frac{3}{\cancel{ 2}}}\implies ? }\)

OpenStudy (anonymous):

8x^3?

OpenStudy (jdoe0001):

close 6 * 3/2 6 divides by 2 and leaves 3 thus \(\bf \cancel{ 6}\cdot \frac{3}{\cancel{ 2}}\implies 3\cdot 3\implies 9\) \(\bf 8x^9\)

OpenStudy (anonymous):

why did you separate the (4) and the (x^6)

OpenStudy (anonymous):

with the exponent

OpenStudy (jdoe0001):

\(\bf (abc)^n\implies a^n\cdot b^n\cdot c^n\)

OpenStudy (anonymous):

what about x^2 + 3x? I know that you have to use f(x+h) - f(x)/h but i don't understand how to put that equation in there

OpenStudy (jdoe0001):

hmmm you that sounds abit cryptic... though is simpler if you post anew offhand I'd say looks like a function translation

OpenStudy (anonymous):

so put it on a new post?